OpenStudy (anonymous):

(1)/(x-2)+(1)/(x+3)=7/6

OpenStudy (e.mccormick):

\(\dfrac{1}{x-2}+\dfrac{1}{x+3}=\dfrac{7}{6}\) So, solve for x. First, you need to remember there is a restriction when things are on the bottomof a fraction. Then you need to work on solving for x.

OpenStudy (anonymous):

@e.mccormick okay how do you get rid of the restriction?

OpenStudy (e.mccormick):

You can't. It just means that solutions that make a fraction invaild (0 on bottom) are not corrrect.

OpenStudy (anonymous):

@e.mccormick So i should get rid denominators ?

OpenStudy (e.mccormick):

Well, you have more than one fraction, so you need a common denominator first.

OpenStudy (anonymous):

@e.mccormick does the common denominator go into the 7/6 as well ?

OpenStudy (e.mccormick):

It can if you want to move it over fisrts, or you can leave that for cross multiplication. Your choice.

OpenStudy (anonymous):

Oh okay thanks for the help @e.mccormick

OpenStudy (e.mccormick):

np. Let me know how it is going.

OpenStudy (e.mccormick):

Looks like a wholw number and a fraction.

OpenStudy (e.mccormick):

I also checked my answer on a graph. Looks really good. Let me know what you get and I can check it. If you have a process problem, I have the details worked out.